OpenStudy (anonymous):

find f(x) and g(x) such that h(x)=(f og)(x) when h(x)=cubed root of x^2-8

OpenStudy (freckles):

hint on how to choose g and f: \[h(x)=\sqrt[3]{x^2-8}=\sqrt[3]{g(x)}=f(g(x))\]